You’re involved in the design of a mission carrying humans to the surface of the 

planet Mars, which has a radius 3.37x106m and a mass of 6.42x1023 kg. The 

earth weight of the Mar’s lander is 39,200 N. Calculate its weight and the 

acceleration due to Mar’s gravity at 6.0 x 106 m above the surface of Mars.


1
Expert's answer
2021-01-23T17:37:58-0500

"W_M=G\frac{mM}{(r+h)^2}=G\frac{WM}{g(r+h)^2}\\
W_M=6.67\cdot10^{-11}\frac{(39200)6.42\cdot10^{23}}{9.8(3.37\cdot10^{6}+6\cdot10^{6})^2}=1951\ N"

"g_M=\frac{1951\cdot 9.8}{39200}=0.49\frac{m}{s^2}"